How many grams of excess reagent are left over after the reaction of 4.36 g of Al and 8.50
g HCl in the following balanced chemical equation ?
1
Expert's answer
2020-03-30T10:58:33-0400

2Al+6HCl=2AlCl3+3H2


We calculate the theoretical mass from the reaction

"m(Al)=n\times M=2\times27=54 g"

"m (HCl)=n\times M=6\times36.5=219 g"

For actual mass HCL 8.50g such actual mass Al is required:


"\frac{8.5}{219}=\frac{x}{54}" , x=2.095


2.095 g is necessary for interaction with 8.50 g of hydrochloric acid

therefore, we find an excess of aluminum reagent

4.36-2.095=2.265